# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v17

@@ Abstract
As a rigorous statistical approach, statistical Taylor expansion extends the conventional Taylor expansion by replacing precise input variables with random variables of known distributions and sample counts to compute the mean, the standard deviation, and the reliable factor of each result. It tracks the propagation of the input uncertainties through intermediate steps, so that the final analytic result becomes path independent. Therefore, it differs fundamentally from common approaches in applied mathematics that optimize computational path for each calculation. Statistical Taylor expansion may standardize numerical computations for analytic expressions. This study also introduces the implementation of statistical Taylor expansion termed variance arithmetic and presents corresponding test results across a wide range of mathematical applications.
Another important conclusion of this study is that numerical errors in library functions can significantly affect results. It is desirable that each value from library functions be accomplished by an uncertainty deviation.
@@ 1 Introduction
Let $x$ and $\delta x$ be a value and its uncertainty deviation, respectively. As an input, $x$ and $\delta x$ are the mean and standard deviation of a measurement [1][2]. If $\delta x=0$, $x$ is a precise value, otherwise, the pair specifies an imprecise value $x\pm\delta x$.
Let $P(x)\equiv\delta x/|x|$ be the statistical precision (hereafter referred to as precision) of $x\pm\delta x$. A smaller $P(x)$ indicate a higher measurement quality of $x\pm\delta x$, and when $P(x)\ll 1$, it is unclear whether $\delta x$ is too poor to measure $x$, or it is a measurement of $x=0$.
Statistical Taylor expansion determines the result $f\pm\delta f$ of a general analytic expression $f(x,\dots)$, and the reliability $\alpha\in[0,1]$ of $f\pm\delta f$ based on the inputs $x\pm\delta x,\dots$ and their corresponding sample counts.
  • Previous studies have examined the effect of input uncertainties on output values for specific cases [3]. Statistical Taylor expansion generalizes these effects as uncertainty bias, as shown in Formula (2.5) and (2.8) in this paper.
  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this paper.
  • Calculating the reliability of $f\pm\delta f$ seems completely new.
Statistical Taylor expansion is superior to all existing numerical arithmetic.
Conventional floating-point arithmetic [6][7][8] can only calculate the result $f$ which contains unknown amount of rounding error [9][10][11]. As a consequence, a floating-point representation with $10^{-16}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision. While statistical Taylor expansion can track rounding errors.
The bounding range in interval arithmetic [12][13][14][15][16] is inconsistent with the statistical nature of an $x\pm\delta x$ pair and it tends to over-estimate the result uncertainty due to its worst-case assumption [17]. On the other hand, statistical Taylor expansion is precise statistically.
Both conventional floating-point arithmetic and interval arithmetic depend strongly on the specific algebraic form of an analytic function, a phenomenon known as the dependency problem [10][11][12][14], which makes the numerical calculations of analytic problems sometimes more like an art than science. In contrast, statistical Taylor expansion is path independent.
To be rigorous in mathematics and statistics, statistical Taylor expansion abandons the significance arithmetic nature of its processor [17].
As a statistical sampling process, stochastic arithmetic [18][19] is computationally expensive, while statistical Taylor expansion provides a direct characterization without sampling.
In this paper:
  1. 1.
    Section 1 compares briefly statistical Taylor expansion with all other numerical arithmetic.
  2. 2.
@@
  4. 4.
    Section 4 lays out standard to validate variance arithmetic.
  5. 5.
    Section 5 illustrates variance arithmetic in computing polynomial, demonstrating its capability in tracing floating-point rounding errors, and its continuity in parameter space.
  6. 6.
    Section 6 applies variance arithmetic to adjugate matrix and matrix inversion, showing that the uncertainty calculation for a linear calculation is no longer linear. It also distinguishes between a distribution test and a value test.
  7. 7.
    Section 7 evaluates variance arithmetic on common mathematical library functions. It also shows that the error deviation should be close to $1$ except near a distributional pole.
  8. 8.
@@
  10. 10.
  11. 11.
    Section 11 concludes with a summary and a discussion.
@@ 2 Statistical Taylor Expansion
2.1 The Uncorrelated Uncertainty Condition
When there is no systematic error [1][2] among the inputs, the uncertainties of the inputs are uncorrelated to each other, even though the inputs may have significant correlations among them. This condition can be characterized quantitatively in statistics as the uncorrelated uncertainty condition [17].
@@ 2.2 Distributional Zero and Distributional Pole
Figure 2: Probability density function for $\tilde{y}=\sqrt{\tilde{x}}$, various values of $\mu$ as indicated in the legend. The variable $\tilde{x}$ follows a Gaussian distribution with the distributional mean $\mu$ and deviation $1$. The horizontal axis is scaled as $\tilde{y}^{2}$.
@@
In both Figure 1 and 2, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, such as $(5\pm 1)^{2}$ in Figure 1 and $\sqrt{5\pm 1}$ in Figure 2, thereby allowing for characterization of the output $f(x)$ using mean $\overline{f(x)}$ and deviation $\delta f(x)$ only.
@@ 2.3 Statistical Taylor Expansion
$\zeta(n)$ $\equiv\int_{\varrho}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.2)
$\zeta(1)$ $=0;$ (2.3)
Let $\tilde{z}\in[\varrho,\kappa]$ where $\varrho,\kappa$ specify the bounding ranges. Formula (2.2) defines the corresponding bound moment $\zeta(n)$, which further satisfies the mean-reverting condition of Formula (2.3) so that $\kappa$ determines $\varrho$. For any symmetric $\rho(-\tilde{z})=\rho(\tilde{z}):\varrho=-\kappa,\zeta(2n+1)=0$. The probability of $\tilde{z}\not\in[\varrho,\kappa]$ is defined as the bounding leakage $\epsilon(\kappa)=1-\zeta(0,\kappa)$.
@@
$\overline{f(x)}$ $=\int_{-\varrho}^{+\kappa}f(x+\tilde{x})\rho(\tilde{x},\mu,\sigma)d\tilde{x}=f(x)+\sum_{n=1}^{\infty}\sigma^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n);$ (2.5)
@@
$=\sum_{n=1}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.6)
An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.4). Formula (2.5) and Formula (2.3) yield the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively. The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the resulting value.
@@
$\overline{f(x,y)}$ $=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$ (2.8)
$\delta^{2}f(x,y)$ $=\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\big(\zeta_{x}(m)\zeta_{y}(n)-\zeta_{x}(i)\zeta_{x}(m-i)\;\zeta_{y}(j)\zeta_{y}(n-j)\big);$ (2.9)
Under the uncorrelated uncertainty condition, Formula (2.8) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.7), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ denote the variance moments for $x$ and $y$, respectively. Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
@@
$\delta^{2}(x\pm y)$ $=\zeta(2,\kappa_{x})(\sigma_{x})^{2}+\zeta(2,\kappa_{y})(\sigma_{y})^{2};$ (2.11)
@@
$\delta^{2}(xy)$ $=\zeta(2,\kappa_{x})(\sigma_{x})^{2}y^{2}+x^{2}\zeta(2,\kappa_{y})(\sigma_{y})^{2}+\zeta(2,\kappa_{x})(\sigma_{x})^{2}\;\zeta(2,\kappa_{y})(\sigma_{y})^{2};$ (2.13)
When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formula (2.10) and (2.11) the convolution results for $x\pm y$ [4], and Formula (2.12) and (2.13) the corresponding results of the product distribution for xyxy [4].
@@ 2.4 One-Dimensional Examples
Formula (2.14) and (2.15) give the mean and variance for $e^{x}$, respectively:
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}\sigma^{n}\zeta(n)\frac{1}{n!};$ (2.14)
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j!\;(n-j)!};$ (2.15)
Formula (2.16) and (2.17) give the mean and variance for $\log(x)$, respectively:
$\overline{\log(x)}$ $=\log(x)+\sum_{n=1}^{+\infty}P(x)^{n}\frac{(-1)^{n+1}\zeta(n)}{n};$ (2.16)
$\delta^{2}\log(x)$ $=\sum_{n=2}^{+\infty}P(x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j(n-j)};$ (2.17)
Formula (2.19) and (2.20) give the mean and variance for $\sin(x)$, respectively:
@@
$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}\sigma^{n}\eta(n,x)\frac{\zeta(n)}{n!};$ (2.19)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.20)
Formula (2.21) and (2.22) give the mean and variance for $x^{c}$, respectively:
$\frac{\overline{x^{c}}}{x^{c}}$ $=1+\sum_{n=1}^{\infty}P(x)^{n}\zeta(n)\begin{pmatrix}c\\ n\end{pmatrix}$; (2.21)
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=2}^{\infty}P(x)^{n}\sum_{j=1}^{n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ n-j\end{pmatrix}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.22)
The input and output in statistical Taylor expansion reflects the inherent characteristics of the calculation, such as $\sigma\rightarrow P(e^{x})$, $P(x)\rightarrow\delta\log(x)$, $\sigma\rightarrow\delta\sin(x)$, and $P(x)\rightarrow P(x^{c})$.
@@ 2.5 Convergence of Variance
$\forall j$ $:\lim_{n\rightarrow+\infty}\zeta(n)-\zeta(j)\zeta(n-j)=\frac{\kappa^{n}}{n};$ (2.25)
@@
[…]
  • Formula (2.22) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.27). which converges when $P(x)\lesssim 1/\kappa$ although the precise upper bound for $P(x)$ varies with $c$.

### v18

@@ Abstract
As a rigorous extension of conventional Taylor expansion, statistical Taylor expansion replaces precise input variables with random input variables of known distributions and sample counts, to compute the mean, standard deviation, and reliability of each result. This approach tracks the propagation of input uncertainties through intermediate steps, making the final result path independent. Therefore, it differs fundamentally from conventional approaches that are usually path dependent, such that it may standardize numerical computations for analytic expressions. This study further presents an implementation of statistical Taylor expansion termed variance arithmetic, and demonstrates its performance across diverse mathematical applications.
The finding of this study also reveals the potential and substantial impact of numerical errors in library functions. The possible link between statistical Taylor expansion and quantum physics is discussed as well.
@@ 1 Introduction
Let $x$ and $\delta x$ denote a value and its uncertainty deviation, respectively. As input, $x$ and $\delta x$ are usually the mean and standard deviation of a measurement, respectively [1][2]. If $\delta x=0$, then $x$ is a precise value; otherwise, the pair specifies an imprecise value $x\pm\delta x$. Let $P(x)\equiv\delta x/|x|$ be the statistical precision (hereafter referred to as precision) of $x\pm\delta x$. A smaller $P(x)$ indicates a higher measurement quality of $x\pm\delta x$.
Statistical Taylor expansion determines the result $f\pm\delta f$ and reliability of a general analytic expression $f(x,\dots)$ on the basis of inputs $x\pm\delta x,\dots$ and their corresponding distributions and sample counts.
  • Previous studies have examined the effect of input uncertainties on output values for specific cases [3]. Statistical Taylor expansion generalizes these effects as uncertainty bias, as shown in Formulas (2.5) and (2.8) of this study.
  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this study.
  • Calculating the reliability $\alpha\in[0,1]$ of $f\pm\delta f$ seems completely new.
As a new approach, statistical Taylor expansion is superior to all existing numerical arithmetic.
Conventional floating-point arithmetic [6][7][8] computes only the result value $f$. Because $f$ contains unknown amount of rounding error [9][10][11], a floating-point representation with $10^{-16}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision. However, statistical Taylor expansion can account for rounding errors as part of the result deviation $\delta f$.
The bounding range in interval arithmetic [12][13][14][15][16] is inconsistent with the statistical nature of an $x\pm\delta x$ pair and it tends to over-estimate the result uncertainty because of its worst-case assumption [17]. Conversely, statistical Taylor expansion is precise statistically.
Both conventional floating-point arithmetic and interval arithmetic depend strongly on the specific algebraic form of an analytic function, a phenomenon known as the dependency problem [10][11][12][14], which can make conventional numerical computation more an art than a science. In contrast, statistical Taylor expansion is path independent.
To ensure mathematical and statistical rigor, statistical Taylor expansion abandons the significance arithmetic nature of its processor [17].
As a statistical sampling process, stochastic arithmetic [18][19] is computationally expensive, whereas statistical Taylor expansion provides a direct characterization without sampling.
The remainder of this paper is organized as follows:
@@
  • Section 4 presents standards for validating variance arithmetic.
  • Section 5 illustrates variance arithmetic in polynomial computation, demonstrating its ability to trace floating-point rounding errors and its continuity in parameter space.
  • Section 6 describes the applications of variance arithmetic to matrix inversion, distinguishing between distribution tests and value tests.
  • Section 7 discusses the evaluation of variance arithmetic on common mathematical library functions, showing the effect of distribution pole.
@@
  • Section 11 concludes with a summary and a discussion of the findings.
@@ 2 Statistical Taylor Expansion
2.1 Uncorrelated Uncertainty Condition
When inputs contain no systematic error [1][2], their uncertainties are uncorrelated, even though their values can have significant correlations. This condition can be characterized quantitatively in statistics as the uncorrelated uncertainty condition [17].
@@ 2.2 Distributional Zero and Distributional Pole
Figure 2: Probability density function for $\tilde{y}=\sqrt{\tilde{x}}$, for various values of $\mu$ as indicated in the legend. The variable $\tilde{x}$ follows a Gaussian distribution with the distributional mean $\mu$ and deviation $1$. The horizontal axis is scaled as $\tilde{y}^{2}$.
@@
In both Figures 1 and 2, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, such as $(5\pm 1)^{2}$ in Figure 1 and $\sqrt{5\pm 1}$ in Figure 2, thereby allowing for characterization of the output $f(x)$ using mean $\overline{f(x)}$ and deviation $\delta f(x)$ only.
@@ 2.3 Statistical Taylor Expansion
$\zeta(n,\kappa)$ $\equiv\int_{\varrho}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.2)
$\zeta(1,\kappa)$ $=0;$ (2.3)
@@
[…]
$\overline{f(x)}$ $=\int_{-\varrho}^{+\kappa}f(x+\tilde{z})\rho(\tilde{z})d\tilde{z}=f(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
@@
$=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
An analytic function $f(x)$ can be accurately evaluated in a range using the Taylor series as shown in Formula (2.4). Formulas (2.5) and (2.3) yield the mean $\overline{f(x)}$ and variance $\delta^{2}f(x)$ of $f(x)$, respectively. The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the resulting value.
@@
$\overline{f(x,y)}$ $=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\delta x)^{m}(\delta y)^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m,\kappa_{x})\zeta_{y}(n,\kappa_{y});$ (2.8)
$\delta^{2}f(x,y)$ $=\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}(\delta x)^{m}(\delta y)^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
$\big(\zeta_{x}(m,\kappa_{x})\zeta_{y}(n,\kappa_{y})-\zeta_{x}(i,\kappa_{x})\zeta_{x}(m-i,\kappa_{x})\;\zeta_{y}(j,\kappa_{y})\zeta_{y}(n-j,\kappa_{y})\big);$ (2.9)
Under the uncorrelated uncertainty condition, Formulas (2.8) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.7), where $\zeta_{x}(m,\kappa_{x})$ and $\zeta_{y}(n,\kappa_{y})$ denote the bound moments for $x$ and $y$, respectively. Although Formula (2.3) is only 2-dimensional problem space, it can be extended easily to any dimension.
@@
$\delta^{2}(x\pm y)$ $=\zeta(2,\kappa_{x})(\delta x)^{2}+\zeta(2,\kappa_{y})(\delta y)^{2};$ (2.11)
@@
$\delta^{2}(xy)$ $=\zeta(2,\kappa_{x})(\delta x)^{2}y^{2}+x^{2}\zeta(2,\kappa_{y})(\delta y)^{2}+\zeta(2,\kappa_{x})(\delta x)^{2}\;\zeta(2,\kappa_{y})(\delta y)^{2};$ (2.13)
When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formulas (2.10) and (2.11) the convolution results for $x\pm y$ [4], and Formulas (2.12) and (2.13) the corresponding results of the product distribution for xyxy [4].
@@ 2.4 One-Dimensional Examples
Formulas (2.14) and (2.15) give the mean and variance for $e^{x}$, respectively:
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\zeta(n,\kappa)\frac{1}{n!};$ (2.14)
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!\;(n-j)!};$ (2.15)
Formulas (2.16) and (2.17) give the mean and variance for $\log(x)$, respectively:
$\overline{\log(x)}$ $=\log(x)+\sum_{n=1}^{+\infty}P(x)^{n}\frac{(-1)^{n+1}\zeta(n,\kappa)}{n};$ (2.16)
$\delta^{2}\log(x)$ $=\sum_{n=2}^{+\infty}P(x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j(n-j)};$ (2.17)
Formulas (2.19) and (2.20) give the mean and variance for $\sin(x)$, respectively:
@@
$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}(\delta x)^{n}\eta(n,x)\frac{\zeta(n,\kappa)}{n!};$ (2.19)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.20)
Formulas (2.21) and (2.22) give the mean and variance for $x^{c}$, respectively:
$\frac{\overline{x^{c}}}{x^{c}}$ $=1+\sum_{n=1}^{\infty}P(x)^{n}\zeta(n,\kappa)\begin{pmatrix}c\\ n\end{pmatrix}$; (2.21)
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=2}^{\infty}P(x)^{n}\sum_{j=1}^{n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ n-j\end{pmatrix}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.22)
The input and output in statistical Taylor expansion reflects the inherent characteristics of the calculation, such as $\delta x\rightarrow P(e^{x})$, $P(x)\rightarrow\delta\log(x)$, $\delta x\rightarrow\delta\sin(x)$, and $P(x)\rightarrow P(x^{c})$.
@@ 2.5 Convergence of Variance
$\forall j$ $:\lim_{n\rightarrow+\infty}\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)=\frac{\kappa^{n}}{n};$ (2.25)
[…]
  • Formula (2.17) for $\log(x\pm\delta x)$ can be approximated by Formula (2.26), which converges when $P(x)<1/\kappa$.
  • Formula (2.22) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.27), which converges when $P(x)\lesssim 1/\kappa$ although the precise upper bound for $P(x)$ varies with $c$.
@@
Statistical Taylor expansion rejects the distributional zero of $\log(x)$ or $(x\pm\delta x)^{c}$ in the range of $P(x)>1/\kappa$ statistically because of the divergence of Formulas (2.26) and (2.27) mathematically, with $\zeta(2n,\kappa)$ providing the connection between these two perspectives.